Please help with this!
sqrt 4x + 1 + 3 = 0
The 4x+1 is inside the radical sign - the +3 = 0 is not.
Thanks

Answer by fractalier(6550)

Okay we have
sqrt(4x + 1) + 3 = 0
sqrt(4x + 1) = -3
now square both sides
4x + 1 = 9
4x = 8
x = 2
now check it...
it doesn't work because the principal root of 9 is regular three...
thus, no solution...